The speed(v) of car in m/s is related to time t in second by the equation: v= 3+12t-3t^ . Find the maximum of the car in km/h?

1
Expert's answer
2022-04-11T16:23:48-0400

"\frac{dv}{dt}=0 \to 12-6t=0 \to t=2."

"v_{max}=v(2)=3+12*2-3*2^2=15 \frac{m}{s}."